Sandra saves 10% of her salary for retirement. This year her salary was $2,000 more than in the previous year, and she saved $4,900. What was her salary in the previous year?
Complete the equation.
0.10(x + ____) = 4,900
Sandra's salary on the previous year was _____.
Please fill in the blanks .. Thank You!!

Answer by Theo(13342)

she saved 4900 which was 10% of her earnings for this year and last.
take 4900 and divide it by .10 and you get 49000 of earnings in the past 2 years.
if you divide 49000 by 2 you get 24,500 for each.
take 1000 off of one and add it to the other and you get 23500 for one and 25500 for the other.
she earned 23500 in the first year and 25500 in the second year.
25500 - 23500 = 2000 difference.
10% of 23500 is 2350
10% of 25500 is 2550
2350 and 2550 = 4900.

algebraically you could solve as follows:
x = amount she earned in the first year.
x + 20000 = amount she earned in the second year.
.10 * x + .10 * (x + 2000) = 4900
factor out the .10 and you get:
.10 * (x + x + 2000) = 4900
simplify by combining like terms to get:
.10 * (2x + 2000) = 4900
divide both sides of this equation by .10 to get:
2x + 2000 = 49000
subtract 2000 from both sides of this equation to get:
2x = 47000
divide both sides of this equation by 2 to get:
x = 23500
that's what she made the first year.
add 2000 to that and you get:
x = 25500
that's what she made the second year.